OpenStudy (anonymous):

Add. 2m+ 4 + m + 2=2m+6 right??

OpenStudy (anonymous):

2m + m = 3m

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

you want to know 2m+ 4 + m + 2= ?

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

So you sum the similar terms. 2m + m and 4 + 2

OpenStudy (anonymous):

2 m + m = 3m and 4 + 2 =6 which gives you 3m + 6